OpenStudy (beccab003):

EASY QUESTION. I will give medal for the best answer. The garden’s length is three feet longer than the width. If the perimeter is 26 feet, choose the best equation for the situation. (w + 3) (w) = 26 4w + 6 = 26 4w + 3 = 26

hartnn (hartnn):

let width be 'w' "length is three feet longer than the width" length in terms of 'w' will be ?

OpenStudy (beccab003):

This is the equation that I thought it should look like: l=(w+3) + w = 26 but that isn't an option.

hartnn (hartnn):

let width be 'w' "length is three feet longer than the width" length = w+3

hartnn (hartnn):

perimeter = 2(width + length)

hartnn (hartnn):

26 = 2(w + (w+3))

hartnn (hartnn):

got this ?

OpenStudy (beccab003):

Yes, I get it.

hartnn (hartnn):

simplify the right side w+w+3 = ... ?

OpenStudy (beccab003):

4w+3=26

hartnn (hartnn):

no

OpenStudy (beccab003):

2w =3 = 26

hartnn (hartnn):

2(w+w+3) =26 2(2w+3) =26 2*2w +2*3 = 26 4w + 6 =26

OpenStudy (beccab003):

OH! I see! you have to do it to both!

OpenStudy (beccab003):

Thanks!

OpenStudy (beccab003):

Can you help me with one more quick one?

OpenStudy (beccab003):

Solve the following system of equations. x + 4y = 3 3x + 12y = 5

hartnn (hartnn):

factor out 3 from the left side of 2nd equation

hartnn (hartnn):

\(3x+12y =3 (..?..)\)

OpenStudy (beccab003):

12=-3x+5

OpenStudy (beccab003):

sorry 12y=-3x+5

hartnn (hartnn):

thats correct, but not required

OpenStudy (beccab003):

So what do you do then?

hartnn (hartnn):

\(3x+12y= 3(x+4y)\) got this ?

OpenStudy (beccab003):

oh yes

OpenStudy (beccab003):

Which equals: 3x + 12y = 3x +12y

OpenStudy (beccab003):

Bytheway the answer options to this question is: x = 3, y = 4 x = 3, y = 12 infinite solutions no solution I think it is no solution-

hartnn (hartnn):

thats correct.

hartnn (hartnn):

see, \(3x+12y = 3(x+4y)\) but x+4y from 1st equation is 3 \(3x+12y = 3(x+4y) = 3\times 3 =9\) but from 2nd equation, 2x+12y =5 so we get 5 =9 which is not possible for any values of x or y hence no solution

OpenStudy (beccab003):

Thanks!!! I understand it a lot better now.

hartnn (hartnn):

welcome :D